# Hao Sun, Weiwei Deng, Feng Sun, Furu Wei, Qi Zhang Microsoft Corporation

### v1

@@ 4.1 Setup
Table 1: Comparison of documents retrieval performance on different datasets, where R@k stands for Recall@k, M@k stands for MAP@k.
SQuAD NQ TriviaQA PAQ RIR
@@
GeAR 0.810 0.874 0.765 0.871 0.515 0.808 0.885 0.965 0.954 0.897
Table 2: Comparison of units localization performance on different datasets, where R@k stands for Recall@k, M@k stands for MAP@k.
(a) Information localization and generation results of GeAR in Question Answer Retrieval scenario.
(b) Information localization and generation results of GeAR in Related Information Retrieval scenario. The sentences in brackets of corresponding colors are the ground truth of the query.
Figure 3: Visualization of information localization of GeAR . In the two scenarios of Question Answer retrieval and Related Information Retrieval, we propose two different queries for one document and highlight the top 10 tokens with the highest cross attention weights for the corresponding queries. The tokens with orange background are for query1, and the tokens with purple background are for query2. We also show the generated results of GeAR.
SQuAD NQ TriviaQA PAQ RIR
@@
61.2 65.3 66.1 61.0 47.4 60.0 88.1 92.4 87.4 87.1
Table 3: Generation performance of GeAR on different tasks.
(a) On QAR tasks.
(b) On RIR tasks.
(c) Localization performance of different layers.
Figure 4: Plots of generation and localization performance on (a) QAR tasks and (b) RIR tasks as training progresses. (c) shown the localization performance at different layers.
4.2 Overall performance
In this section, we present the overall performance on Documents Retrieval, Units Localization, and Information Generation.
Documents Retrieval Firstly, we report the comparison with existing methods on documents retrieval task in Table 1. The results demonstrate that GeAR delivers competitive performance across multiple datasets, even with limited training data. As a reference, the pre-trained SBERT model used 1.17B sentence pairs, with partial overlap between its training data and our evaluation data. To ensure a fair comparison, we retrained SBERT [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2. and BGE [^3] https://github.com/FlagOpen/FlagEmbedding. using their open source training pipelines, aligned training data and initialization settings. As shown in the retrained model section in Table 1, GeAR achieves superior performance, underscoring the effectiveness of our training approach.
Units Localization Next, we evaluate the performance of each method on the units localization task. In the evaluation process, we provide the query and the document $(q,d)$ to the model and observe whether it is able to find the corresponding fine-grained unit $u$. For the retrieval model, we split the documents into sentences and compute their similarity to the query independently, selecting the top-k sentences. In contrast, GeARlocates units based on the cross attention weights for each sentence given the document and the query, as described in Section 3.5. The results are reported in Table 2. We found that GeAR came out ahead on all metrics, and that GeAR did not require further chunking and encoding of the document. It is observed that SBERTRT and BGERT exhibit suboptimal performance, as their training objective focus solely on optimizing the overall semantic similarity between the document and the query, neglecting the fine-grained semantic relationships. In contrast, GeAR benefits from the joint end-to-end training of retrieval and generation tasks, enabling it not only to retrieve documents closely aligned with the query but also to effectively attend to fine-grained information within the document.
Information Generation Although generation serves only as an auxiliary task in GeAR, we are nonetheless interested in evaluating its generation performance. Table 3 reports the Exact Match (EM) and F1 scores on the QA datasets, and the Rouge Lin (2004) scores on the RIR dataset. Notably, GeAR achieves strong performance on test sets with distributions similar to the training data, such as PAQ and RIR, and performs reasonably well on other test sets. Additionally, Figure 3 illustrates examples of GeAR’s ability to generate correct answers and relevant information, demonstrating its satisfactory generation capabilities.
@@ 4.3 Analysis
Visualization of Information Localization Figure 3 illustrates the information localization and generation results of GeAR across different scenarios. We provide two distinct queries for one document and highlight the top 10 tokens with the highest cross attnetion weights corresponding to each queries. In Figure 3(a), the two queries focus on time and location, respectively. GeAR not only gave the correct answers to both queries but also dynamically adjusts its query-specific focus: it assigns higher attention weights to time-related tokens in response to the first query and prioritizes tokens related to countries and regions in response to the second query. In Figure 3(b), GeAR will focus on the definition of the AVL tree itself, and the insertion, deletion, rotation and rebalancing of the AVL tree, and generate corresponding sentences. It can be seen that the added generation task has brought improvements to the model in terms of performance and qualitative effects, making it accurate in localization and generation.
Correlation of Generation and Localization In this section, we analyze the relationship between the generation and localization tasks. As illustrated in Figure 4(a) and 4(b), we plot the performance coordinates from epoch 1 to epoch 10 during training, where the horizontal axis represents the generation performance and the vertical axis represents the localization performance. The results reveal a strong correlation between the two tasks. This observation demonstrates that the generation task, designed as a proxy, effectively enhances the model’s ability to extract fine-grained information from documents. These findings highlight the synergistic relationship between generation and localization.
@@
Localization performance of different layers In GeAR, the fusion encoder and decoder interact through the cross attention module at each layer. To investigate the relationship between localization performance and model depth, we plot the localization performance using cross attention weights across different layers in Figure 4(c). The results indicate that high-level token embeddings perform well, as they capture rich semantic information through deeper layers of the network. Interestingly, we observe that the highest layer does not yield the best localization performance. Instead, peak performance is achieved in the last 3 to 4 layers [^4] In this work, we utilized the 10th layer for evaluation.. This phenomenon may arise because the representations in the highest layer are optimized to serve the final task rather than intermediate localization. Similar findings have been reported in prior studies involving encoder-only and decoder-only models Jawahar et al. (2019); Skean et al. (2024).
The Affect of Language Modeling Objectives
In this work, we utilize only the information corresponding to the query as supervision and incorporate a language modeling objective. It enables the model to achieve stronger capabilities in both information localization and generation, without requiring additional loss functions or complex module designs. However, as a trade-off, we observe a slight decrease in retrieval performance when compared to using only the contrastive learning objective for the retrieval task, as shown in the last two rows of Table 1. How to further design the balance between the two training objectives from the perspective of multi-task learning so that they benefit from each other is a point that can still be explored in the future.
@@ 5 Conclusion
In this work, to address the challenges of unexplainable and coarse-grained results inherent in current bi-encoder retrieval methods, we propose a direct and effective modeling method: Generation Augmented Retrieval (GeAR). GeAR enhances fine-grained information localization and generation capabilities by incorporating a decoder and a lightweight cross-attention layer, while maintaining the efficiency of a bi-encoder. Experimental results across multiple retrieval tasks and two different scenarios demonstrate that GeAR achieves competitive performance. Furthermore, its ability to accurately and reasonably localize information makes it particularly promising for downstream tasks such as web search, semantic understanding, and retrieval-augmented generation (RAG). We hope this work offers valuable insights into the gradual unification of natural language understanding and generation paradigms, paving the way for more versatile and explainable retrieval systems in the future.
@@ Appendix A Data Construction
For the construction of Relevant Information Retrieval data, we have also tried to collect paired sentences and make LLM expand one of them into a document. However, we found that other sentences in the LLM expansion were less informative than the original sentence, for example, being some descriptive statements were generated around the original sentence. This pattern tends to cause the model to learn to locate the central sentence, or the most informative sentence, in the expanded document, leading model to ignore the query. So please be aware of this if you plan to try this way of constructing your data.
@@ B.1 Training
Table 6: The evaluation data statistics for the document retrieval task.
@@
Table 7: The evaluation data statistics for the units localization and information generation tasks.
@@ B.2 Evaluation
Documents Retrieval First, for the document retrieval task, the queries come from the test set in the respective dataset, and the candidate documents are all documents within the entirety of the dataset, including the SQuAD (Rajpurkar et al., 2016), NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and RIR datasets. It is difficult to encode all the documents of the PAQ dataset because the dataset is too large. So for the PAQ dataset, we sampled 1M documents and 20k queries, all of which have no intersection with the training data. The evaluation data statistics for the document retrieval task are shown in Table 6.
Units Localization and Information Generation  For these two tasks, we directly use the test set data corresponding to the respective datasets. Therefore, their number is consistent with the number of queries in Table 6. For the RIR dataset, we sample 10k records as the test set. The evaluation data statistics for the units localization and information generation tasks are shown in Table 7.
@@ Appendix C HyperParameters and Implementation Details
Appendix D More Visualization

### v2

@@ 4.2 Overall performance
Table 1: Comparison of global documents retrieval performance on different datasets, where R@k stands for Recall@k, M@k stands for MAP@k.
Method SQuAD NQ TriviaQA PAQ RIR
@@
Reranker model
BGE-Reranker-B 0.690 0.749 0.641 0.740 0.399 0.640 0.690 0.762 0.884 0.850
BGE-Reranker-L 0.751 0.813 0.670 0.770 0.464 0.737 0.704 0.778 0.891 0.873
GeAR 0.814 0.878 0.761 0.865 0.510 0.797 0.884 0.965 0.933 0.897
[…]
Table 2: Comparison of local information retrieval performance on different datasets, where R@k stands for Recall@k, M@k stands for MAP@k.
(a) Local information retrieval and generation results of GeAR in Question Answer Retrieval scenario.
(b) Local information retrieval and generation results of GeAR in Related Information Retrieval scenario. The sentences in brackets of corresponding colors are the ground truth of the query.
Figure 3: Visualization of local information retrieval of GeAR . In the two scenarios, we pose two different queries for each document and highlight the top 10 tokens with the highest cross attention weights. The tokens with orange background are for query1, with purple background are for query2. We also show the generated results of GeAR.
Local Information Retrieval Next, we evaluate the performance of each method on the local information retrieval task. In the evaluation process, we provide the query and the document $(q,d)$ to the model and observe whether it is able to retrieve the corresponding fine-grained unit $u$. For the retrieval model, we split the documents into sentences and compute their similarity to the query independently, selecting the top-k sentences. In contrast, GeAR retrieves units based on the cross attention weights for each sentence given the query, as described in Section 3.5. The results are reported in Table 2.
It is observed that SBERTRT and BGERT perform mediocrely, as their training objective focus solely on optimizing the overall similarity between the document and the query, neglecting the fine-grained semantic relationships. The more complex BGE-reranker model performs better than the pure retrieval model. GeAR leads the way in all metrics, showing an average relative improvement of 12.9% over the suboptimal BGE-Reranker-L. Notably, GeAR does not require further chunking and encoding of the document. In contrast, GeAR benefits from the joint end-to-end training of retrieval and generation, enabling it not only retrieve documents closely aligned with the query but also effectively retrieve fine-grained information within the document.
Method SQuAD NQ TriviaQA PAQ RIR
@@
Llama 3.2 3B 60.7 73.3 57.7 59.9 50.4 66.7 62.7 75.5 69.4 67.9
Llama 3.3 70B 66.2 77.7 61.0 66.9 56.6 73.0 61.0 74.5 84.4 84.0
GeAR 60.0 64.5 65.7 60.7 46.5 59.1 87.5 91.9 87.6 87.3
Table 3: Generation performance on different datasets.
@@ 4.3 Analysis
$\alpha$ Global Retrieval Local Retrieval
Ave Recall Ave Recall
0 0.844 0.663
0.25 0.846 0.781
0.5 0.844 0.785
0.75 0.839 0.784
1 0.838 0.784
Table 4: Comparison of performance on two retrieval tasks when the LM loss weight $\alpha$ is varied.
The Effect of Language Modeling Objectives In this work, we not only optimize the retrieval performance through contrastive learning, but also enhance GeAR through the information generation task of a given query, so that it has fine-grained semantic understanding and content retrieval capabilities. We find that if LM loss is removed, both global and local retrieval performance of the model is reduced, as shown in the last row of Table 2 and Table 2. Further, we also explore the impact of the weight of LM loss on the overall performance. In Table 4, we observed that the effect of the generation on the retrieval performance is inverted U-shaped, with the optimal values at 0.25 and 0.5 respectively. Higher weights may cause the model to focus on learning the generation task instead, which is similar to previous findings (Sener and Koltun, 2018).
@@
Visualization of Local Information Retrieval The key distinction between GeAR and traditional retriever is its ability to mine the local information within the document that is most relevant to the query. Figure 3 illustrates this process and the generation results of GeAR across different scenarios. For each document, we provide two distinct queries and highlight the top 10 tokens with the highest cross attnetion weights corresponding to each query. In Figure 3(a), the two queries are related to time and location respectively. GeAR not only provides the correct answers but also dynamically adjusts its query-specific focus: it assigns higher attention weights to time-related tokens to the first query and prioritizes tokens related to countries and regions to the second query. In Figure 3(b), depending on the query, GeAR focuses on the concept of AVL tree, as well as operations such as insertion and rebalancing, generating corresponding sentences. It is evident that the added generation task enhances the accuracy of local information retrieval. Furthermore, GeAR can not only retrieve detailed content related to the query but also generates corresponding text for reference. This advancement shifts the retrieval results from being mere numerical values to more intuitive and explainable.
Figure 4: Local information retrieval performance of different layers.
@@
Local Retrieval Performance of Different Layers In GeAR, the query and document tokens interact through the cross attention module at each layer of the fusion encoder. In Figure 4, we plot the local retrieval performance using cross attention weights across different layers to examine its relationship with model depth. The results indicate that higher layers generally perform well, as the token embeddings at these layers capture rich semantic information. Interestingly, we observe that the highest layer does not yield the best performance. Instead, peak performance is reached in the last 3 to 4 layers [^4] In this work, we utilized the 10th layer.. This phenomenon may arise due to the representations in the highest layer are optimized to serve the final task rather than intermediate interactions. Similar observations have been reported in previous studies involving encoder-only and decoder-only models Jawahar et al. (2019); Skean et al. (2024).
Information Generation Although generation serves as an auxiliary task in GeAR and the decoder is lightweight, we are nonetheless interested in its generation performance. Table 3 reports the Exact Match (EM) and F1 scores on the QA datasets, and the Rouge Lin (2004) scores on the RIR dataset. For reference, we include results from the Llama series model (Dubey et al., 2024). Notably, GeAR achieves surprising performance on the in-domain data, and performs reasonably well on other test sets. Additionally, Figure 3 illustrates examples of GeAR’s ability to generate answers and relevant information, showcasing its satisfactory generation capabilities.
@@ 5 Conclusion
In this work, to address the challenges of unexplainable and coarse-grained results inherent in current bi-encoder retrieval methods, we propose a direct and effective modeling method: Generation Augmented Retrieval (GeAR). GeAR enhances fine-grained information retrieval by introducing a generation task and incorporating a lightweight decoder and cross attention module, while maintaining the efficiency of the bi-encoder. Experimental results across multiple retrieval tasks and two different scenarios demonstrate that GeAR achieves excellent performance and have both global and local understanding and retrieval capabilities. Qualitative analysis further highlights its intuitive and explainable retrieval results. These capabilities make GeAR particularly promising in downstream tasks such as web search and retrieval-augmented generation (RAG). We hope that this work offers valuable insights into the gradual unification of natural language understanding and generation paradigms, paving the way for more general and explainable retrieval systems in the future.
@@
Thirdly, the decoder of GeAR has only 110M parameters, the same as the encoder. Moreover, the focus of GeAR is not to optimize the generation performance of the model, and the generation task is not the main task. Therefore, GeAR cannot complete other complex generation tasks like Llama (Dubey et al., 2024). In future work, whether GeAR can be scaled up to enable it to complete retrieval tasks and respond well to various generation problems will be an interesting direction.
@@ Appendix A Data Construction
For the construction of Relevant Information Retrieval data, we have also tried to collect paired sentences and make LLM expand one of them into a document. However, we fine that other sentences in the LLM expansion were less informative than the original sentence, for example, being some descriptive statements were generated around the original sentence. This pattern tends to cause the model to learn to locate the central sentence, or the most informative sentence, in the expanded document, leading model to ignore the query. So please be aware of this if you plan to try this way of constructing your data.
@@ B.1 Training
Table 7: The evaluation data statistics for the global document retrieval task.
@@
Table 8: The evaluation data statistics for the local information retrieval and generation tasks.
@@ B.2 Evaluation
Global Documents Retrieval First, for the global document retrieval task, the queries come from the test set in the respective dataset, and the candidate documents are all documents within the entirety of the dataset, including the SQuAD (Rajpurkar et al., 2016), NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and RIR datasets. It is difficult to encode all the documents of the PAQ dataset because the dataset is too large. So for the PAQ dataset, we sampled 1M documents and 20k queries, all of which have no intersection with the training data. The evaluation data statistics for the document retrieval task are shown in Table 7.
Local Information Retrieval and Generation  For these two tasks, we directly use the test set data corresponding to the respective datasets. Therefore, their number is consistent with the number of queries in Table 7. For the RIR dataset, we sample 10k records as the test set. The evaluation data statistics for the local information retrieval and generation tasks are shown in Table 8.
@@ Appendix C HyperParameters and Implementation Details
CPU Ratio (vs BGE) GPU Ratio (vs BGE)
BGE 225s 1x 34s (1 * H20) 1x
GeAR 288s 1.28x 56s (1 * H20) 1.65x
BGE Reranker-large - - 285s (8 * H20) 8.38x (8 * H20)
Table 9: Comparison of the inference cost of the models on the local information retrieval task.
Appendix D Inference Cost
In our method, GeAR implements three distinct forward processes:
Global Retrieval: Uses only the bi-encoders for global retrieval, with computational complexity identical to classical retrieval models. Document embeddings can be precomputed offline.
Local Retrieval: Computes fusion encoder cross-attention weights for local retrieval without decoder involvement.
Generation: Activates the decoder only when needed, to generate text.
In the local retrieval, GeAR may introduce some inference cost compared to the classic bi-encoder. Therefore, we tested the inference cost on 2889 test data of NQ task in a CPU environment (AMD EPYC 9K84 96-Core Processor * 2) and a GPU (NVIDIA H20) environment, the results are reported in Table 9.
We observe that GeAR consumes about 1.28x (on CPU) and 1.65x (on GPU) of the same-sized bi-encoder, which is a moderate constant time increase. We also tested the inference speed of BGE Reranker-large on 8 * H20. Since BGE reranker is a complex cross encoder, it needs to fully interact and score each query and candidate. Therefore, its time consumption on 8 * H20 is still 8.38x that of BGE.
Appendix E More Visualization